jamie made a veggie tray for her friend's birthday party. It had celery $1.25/stalk and corrots cost $.95/bunch. It was a big party, so sher purchased 18 veggies and paid $20.40. Write a system of equations and solve it.
Answer by Cromlix(4381)

Hi there,
Make celery = x
Make carrots = y
x + y = 18.......(1)
125x + 95x = 2040.......(20
Multiply Eq (1) by 125
125x + 125y = 2250......(1)
125x + 95y = 2040.......(2)
Subtract Eq (2) from (1)
......30y = 210
......y = 210/30
......y = 7
Substitute y = 7 into Equation (1)
x + y = 18
x + 7 = 18
x = 18 - 7
x = 11
Jamie bought 11 stalks of celery
and 7 bunches of carrots.
Hope this helps :-)